OpenStudy (anonymous):

a little help over here? find the 9th term of the sequence described by: A(n)=7+(n-1)(3) A) 42 B) 9 C) 31 D) 27

OpenStudy (anonymous):

What grade u in?

OpenStudy (anonymous):

A(9) = 7 + (9-1)(3)

OpenStudy (anonymous):

9

OpenStudy (anonymous):

OOOOOOOOOOOO

OpenStudy (templeguy):

D

OpenStudy (anonymous):

6

OpenStudy (anonymous):

A(9) = 7 + (8)(3) =7+24 =31

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

c^

OpenStudy (anonymous):

The answer is NOT D it is C.

OpenStudy (anonymous):

thx ^_^